ICM 1998, Berlin, Aug. 18–27

Abstracts of Plenary and Invited Lectures

Section:

1. Logic

1991 MS Classification: 3, 26, 51 Wilkie, Alex J., Oxford University, UK O-minimality

The study of 0-minimal structures, also known as geometric categories, is a branch of model theory but it has potential geometrical interest. The aim of the pure theory, which I shall review briefly, is to generalise many of hereditary, geometrical and finiteness properties enjoyed by the category of subanalytic sets and maps (eg Whitney stratification, triangulation, Lojasevic inequality) to a general axiomatic setting. The second half of the talk will focus on examples. In particular I shall indicate why (real) functions satisfying Pfaffian differential equations live in O-minimal structures.

ICM 1998 Berlin · Plenary and Invited Lectures